OpenStudy (idealist10):

Solve the initial-value problem y'-xy=xy^(3/2), y(1)=4.

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (anonymous):

\[\begin{align*} y'-xy&=xy^{3/2}\\ y'&=xy+xy^{3/2}\\ y'&=xy(1+y^{1/2})\\ \frac{dy}{y(1+\sqrt y)}&=x~dx \end{align*}\] Is this what you have so far?

OpenStudy (idealist10):

I don't have that so far because I did the Bernoulli method. So can I do it the way you did instead of Bernoulli?

OpenStudy (anonymous):

Yes, the equation is separable. I'm not sure the Bernoulli substitution works for rational values of \(n\). Let's check... \[\begin{align*} y'-xy&=xy^{3/2}\\ y^{-3/2}y'-xy^{-1/2}&=x \end{align*}\] So we'd use the substitution, \(v=y^{-1/2}\), which gives \(v'=-\dfrac{1}{2}y^{-3/2}y'\), so \[\begin{align*} 2v'-xv&=x\\ v'-\frac{x}{2}v&=\frac{x}{2} \end{align*}\] I guess it's still valid! Find your integrating factor and you'll be set.

OpenStudy (idealist10):

I got v'-x/2*v=-x/2 instead of x/2 like yours. And the integrating factor I got is e^(x^2/4) but how do I integrate (x/2)(e^(x^2/4))?

OpenStudy (anonymous):

Oh yeah that's a typo on my part, sorry. Should be \[\begin{align*} \color{red}{-2}v'-xv&=x\\ v'+\frac{x}{2}v&=-\frac{x}{2} \end{align*}\]

OpenStudy (idealist10):

So do I do integration by parts to integrate (-x/2)(e^(x^2/4))?

OpenStudy (anonymous):

\[\mu(x)=\exp\left(\int\frac{x}{2}~dx\right)=\exp\left(\frac{1}{4}x^2\right)=e^{(x^2)/4}\] \[\begin{align*} e^{(x^2)/4}v'+e^{(x^2)/4}\frac{x}{2}v&=-e^{(x^2)/4}\frac{x}{2}\\ \frac{d}{dx}\left[e^{(x^2)/4}v\right]&=-e^{(x^2)/4}\frac{x}{2}\\ e^{(x^2)/4}v&=-\frac{1}{2}\int xe^{(x^2)/4}~dx \end{align*}\] IBP isn't necessary here. Just substitute \(t=\dfrac{x^2}{4}\), so \(dt=\dfrac{x}{2}~dx\), or \(2~dt=x~dx\). \[\begin{align*}e^{(x^2)/4}v&=-\int e^t~dt\\ e^{(x^2)/4}v&=-e^t+C\\ e^{(x^2)/4}v&=-e^{(x^2)/4}+C\\ v&=-1+Ce^{-(x^2)/4} \end{align*}\]

OpenStudy (idealist10):

Let me work it out.

OpenStudy (idealist10):

So y=(-1+(e^(x^2/4))/C)^2, right?

OpenStudy (anonymous):

We used \(v=y^{-1/2}\), which means \(y=\dfrac{1}{v^2}\). You should then have \[y=\frac{1}{\left(Ce^{-(x^2)/4}-1\right)^2}\]

OpenStudy (idealist10):

Got it! Thanks!

OpenStudy (anonymous):

You're welcome!

OpenStudy (idealist10):

:)